# Redefining Neural Operators in $d+1$ Dimensions

### v2

@@ Appendix C Experiment Settings and Results
    where $\rho=\rho(x,t)$ and $\bm{u}=\bm{u}(x,t)$ denote the density and the velocity vector field, respectively. $p$ is the pressure, $\bm{\tau}=\bm{\tau}(\bm{u},\rho;\nu)$ is the eviatoric stress tensor with kinematic viscosity $\nu$, $\bm{g}$ represents the gravity, and $\kappa$ is the coefficient of molecular diffusivity. A key dimensionless parameter in the RT simulations is the Atwood number $A=(\rho_{h}-\rho_{l})/(\rho_{h}+\rho_{l})$, which quantifies the density contrast between the two fluids. For example, miscible fluids with greater density contrasts exhibit stronger instabilities. In this work, we adopt the dataset with $A=0.25$ provided in (Ohana et al. (2024)) to train a solution operator $vec(\rho,\bm{u};t=t_{0})\mapsto vec(\rho,\bm{u};t=t_{0}+20)$ which evolves the system from rest states (with small perturbations) to mixed states after 20 time steps. The models are trained on 220 mapping pairs and tested on 40 pairs from distinct trajectories, using a downsampled grid of size $32\times 32\times 32$. Horizontal periodic boundary conditions and impermeable vertical walls are imposed. See Tab.7 for results.
@@
    To compare the baseline models in a global context of spatio-temporal dynamics, we evaluate their performance on predicting the next-hour ERA5 global wind field (Hersbach et al. (2020)). All models are trained using data from the previous four days to predict the next-hour wind field at a grid size of $512\times 512$, and are tested on data from a new day with the same grid size. See Tab.20 for results.
We emphasize the broader learning difficulty spectrum of the benchmarks we selected. As shown in Fig. 5 (Li et al. (2021)) which illustrates the spectral decay of truncation modes, our chosen benchmarks 6 and 7 contain significantly more high-frequency modes than simpler benchmarks such as 3 or other incompressible NS benchmarks with $\nu>1\times 10^{-4}$. To further evaluate the capability of these efficient and lightweight neural operators in capturing complex non-linear dynamics in 3D space, we include two additional, extremely challenging benchmarks: 8 and 9. Moreover, we also compare the performance of all baseline models in the rapid prediction of global wind fields, providing a comprehensive evaluation across both highly nonlinear and large-scale spatio-temporal dynamics.
@@
  • Mainly two types of Positional Features (Pos. Feat.), absolute coordinates (Abs.) and reference grid (Ref.), are concatenated into input signals: (1) FNO (Li et al. (2021)) concatenates the input function value $a\in\mathbb{R}^{d_{a}}$ with the normalized absolute position coordinate $x\in\mathbb{R}^{d}$, usually scaled between 0 and 1. (2) Transolver (Wu et al. (2024)), on the other hand, uses an advanced version which constructs a coarse reference grid and computes Euclidean distances from each input point to each point in this reference grid, concatenating these relative positions $x_{ref}\in\mathbb{R}^{(d_{ref})^{d}}$. In our work, we utilize the same absolute positional features as FNO to ensure a fair comparison.
@@
  • For Transolver, the optimizers and scheduler type is AdamW (Loshchilov & Hutter (2019)) and OneCycleLR, while for FNO and SKNO are Adam (Kingma & Lei Ba (2015)) and Cosine Annealing. All the learning rates are initially set to $0.001$. More configurations can be found in the code provided in our supplementary materials.
@@
Appendix D Related Work
@@
Unlike traditional methods (Lucia et al. (2004); Chinesta et al. (2011)) that rely on dimension reduction techniques, such as symmetry analysis (Cantwell (2002)), to represent equations within their original domain, approaches like the level set method (Osher & Fedkiw (2001); Osher et al. (2004); Adalsteinsson & Sethian (1995); Jin & Osher (2003)) and the recent Schrödingerisation method (Jin & Liu (2024); Jin et al. (2023; 2024)) propose extending original PDEs to one more higher dimensional domains to create a pipe for functions to evolve. Similar "adding one dimension" ideas have also been used in signal processing like wavelets (Meyer (1992)) which employ scale and shift dimensions to represent original signals, and in quantum mechanics like the Wigner functions (Wigner (1932)) which utilize a single function displaying the probability (quasi-)distribution in the phase space.
@@
In scenarios where PDEs are unknown or training data is limited, supervised learning methods are employed to approximate solution operators, which recently shows their capacity to universally represent solution operators for partial differential equations (PDEs) (Hornik et al. (1989); Kovachki et al. (2021); Lu et al. (2021); Lagaris et al. (1998)). Existing approaches mainly fall into two paradigms: Common Neural Networks and Neural Operators.
@@
Common NNs approximate solution operators working on a fixed grid size, which maps discretized input to output signals using architectures such as MLPs (He et al. (2016); Chen et al. (2018)), CNNs (Qu et al. (2022); Brandstetter et al. (2023)). For example, (Obiols-Sales et al. (2020); Khoo et al. (2021)) utilizes a typical CNN structure to map the input to the output image, which is composed of the solution of numerical solvers.
@@
Neural operators (NOs) are designed to map between function spaces, enabling generalization across varying input discretizations without retraining (Kovachki et al. (2023); Wang et al. (2023)). One of the most popular NO definitions is the kernel integral operator (Li et al. (2021); Raonic et al. (2023b); Cao et al. (2024)). Early kernel function implementations employed graph-based message-passing networks (Li et al. (2020a; b)), where dense edge connections encoded these kernels. However, these methods faced scalability challenges due to long training times and inefficiencies in modeling interactions across multiple spatial and temporal scales. Transformer-based models (Kissas et al. (2022)), similar to graphs, also encounter limitations in computational efficiency due to quadratic time complexity. To address these challenges, tricky computation strategies on kernels have been proposed. Galerkin-based methods (Cao (2021); Ern & Guermond (2004)) employ linear projections to parameterize kernels, while (Xiao et al. (2024)) introduces a decomposition constrained by positive semi-definite assumptions. However, such methods suffer from the problems of reduced expressiveness due to the absence of non-linear activations like softmax. Recent SOTA model transolver (Wu et al. (2024)) proposes a physics-attention module to approximate the kernel function with the aid of a self-attention-based aggregation on flattened data. While these graph or transformer-based advanced methods try to escape from the heavy quadratic complexity by tricky implementations on the kernel functions originally defined in (Li et al. (2020a; 2021)), they still pay the computational price for their neglect of the framework and model alignment in embedding spaces.
@@
Additionally, Fourier Neural Operators (FNO) (Li et al. (2021)) efficiently implement spectral kernel representations by global convolutions via the Fast Fourier Transform (FFT). While training fast, it also faces the representation bottleneck like other NO methods because of the unaligned evolution in embedding spaces.
@@
There is another NO definition: DeepONet (Lu et al. (2021)) employs a branch-trunk architecture inspired by the universal operator approximation theorem (Chen & Chen (1995)). These representations are combined through a dot product for the embeddings composed with functional outputs to predict the solution values. Recent advancements, such as integrating numerical solvers with DeepONet (Zhang et al. (2024)), have improved convergence by capturing multi-frequency components, enhancing performance on complex PDE systems. However, its general-purpose design may limit its effectiveness in scenarios requiring high accuracy and PDE-aligned structural representations.
Appendix E Properties of SKNO
Complexity Analysis
@@ Appendix E Properties of SKNO
To analyze computational complexity across general dimensional settings (1D, 2D, 3D, …), let $N_{x}$ denote the number of grid points in the spatial domain, $N_{p}$ the grid size along the auxiliary dimension, and $k$ the number of aggregated slices/nodes (used in Transolver). The complexity comparison is summarized in Tab. 16.
Table 16: Complexity comparison of different models.
@@
Transolver $\mathcal{O}\!\left[N_{p}(N_{x}k+k^{2})\right]\;(k\geq 32,\;N_{p}=128\text{ or }256)$
FNO $\mathcal{O}\!\left[N_{p}(N_{x}\log(N_{x}))\right]\;(N_{p}\leq 32)$
SKNO (Ours) $\mathcal{O}\!\left[N_{p}(N_{x}\log(N_{x}\cdot N_{p}))\right]\;(N_{p}\leq 32)$
[…]
In lower dimensions like 1D or 2D cases, where $N_{x}$ is exponentially smaller, our method retains the linear-complexity advantages while achieving superior performance with significantly smaller scale constants compared to the SOTA Transolver (under the log benefit from Fast Fourier Transform), which we have validated across various benchmarks (see Tab.7, Tab.18 and 20).
Spectral Interpretation
Since our SKNO adopts a spectral implementation, we primarily compare it to the FNO (Li et al. (2021)) as a reference framework. This choice also facilitates generalization to other spectral models like (Guibas et al. (2022); Li et al. (2024a)). To avoid notational clutter, we begin with the $d=1$ case, noting that the arguments extend straightforwardly to higher dimensions.
Theorem 1 (Neural Operator along the Auxiliary Dimension).
@@
$\hat{\bm{v}}_{i}:=F^{x}_{k_{N}}[P^{x}_{k_{N}}v](\hat{\bm{x}}_{i},\bm{p})\in\hat{V}_{i}\subset\mathbb{C}^{d_{v}},\qquad\tilde{\hat{\bm{v}}}_{i}:=F^{p}[R^{p}_{k_{M}}(F^{x}_{k_{N}}[P^{x}_{k_{N}}v]](\hat{\bm{x}}_{i},\tilde{\bm{p}}))\in\tilde{\hat{V}}_{i}\subset\mathbb{C}^{d_{v}},$
@@
    Approximating an operator $\mathcal{M}$ is equivalent (via the Fourier-conjugate decomposition) to approximating the conjugate mapping(s) $\{\hat{\mathcal{M}}_{i}\}_{i=1}^{k_{N}}$ (and, after also Fourier-transforming in $p$, the conjugates $\{\tilde{\hat{\mathcal{M}}}_{i}\}$).
@@
    The forward DFT/frequency-projection and the inverse discrete transform admit neural operator approximation on the spectral domain (See Lemma 7 and Lemma 8 in (Kovachki et al. (2021))), hence the operator learning problem reduces to a finite-dimensional neural-network approximation on the compact coefficient sets $\hat{V}_{i},\ \tilde{\hat{V}}_{i}$.
@@
Note. $\hat{\mathcal{M}}_{i}:\hat{V}_{i}\subset\mathbb{R}^{2d_{v}}\rightarrow\mathbb{R}^{2d_{v}}$ for $i=1,...,k_{N}$ and $\tilde{\hat{\mathcal{M}}}_{i}:\tilde{\hat{V}}_{i}\subset\mathbb{R}^{2d_{v}}\rightarrow\mathbb{R}^{2d_{v}}$. Usually, we take $k_{M}=d_{v}$ and the $\frac{length(\bm{p})-k_{M}}{length(\bm{p})}$ random dropout $R^{p}_{k_{M}}=I^{p}$.
Proof. By the Fourier-conjugate decomposition (Eq. (12) in (Kovachki et al. (2021))), the operator can be written as
$\mathcal{M}$ $=\;(F^{x}_{k_{N}})^{-1}\circ\hat{\mathcal{M}}_{k_{N}}\circ F^{x}_{k_{N}}\circ P^{x}_{k_{N}}$
$=\;(F^{x}_{k_{N}})^{-1}\circ(F^{p})^{-1}\circ\tilde{\hat{\mathcal{M}}}_{k_{N}}\circ F^{p}\circ F^{x}_{k_{N}}\circ P^{x}_{k_{N}}.$ (33)
@@
The mapping in Equ.33 $F^{p}F^{x}_{k_{N}}[P^{x}_{k_{N}}(\cdot)]$ sends the compact input family $V$ into a compact subset of Fourier coefficient space. Lemma 7 and Lemma 8 of (Kovachki et al. (2021)) guarantee that the left-hand and right-hand parts of both $\hat{\mathcal{M}}_{k_{N}}$ and $\tilde{\hat{\mathcal{M}}}_{k_{N}}$ can be approximated within the desired accuracy.
@@
Thus, the problem reduces to approximating the conjugate operators $\hat{\mathcal{M}}_{k_{N}}$ or $\tilde{\hat{\mathcal{M}}}_{k_{N}}$, both of which are continuous maps defined on compact subsets of $\mathbb{R}^{2d_{v}}$. By the universal approximation theorem for neural networks (Barron (2002); Hornik et al. (1989)), any such continuous map can be approximated to arbitrary accuracy by a finite neural network. Hence, there exists finite NNs that approximate the conjugate mappings on $\hat{V}_{i}$ (or $\tilde{\hat{V}}_{i}$) within the desired accuracy.

### v3

@@ Appendix B Model, Experiment Settings and Results
    where $\rho=\rho(x,t)$ and $\bm{u}=\bm{u}(x,t)$ denote the density and the velocity vector field, respectively. $p$ is the pressure, $\bm{\tau}=\bm{\tau}(\bm{u},\rho;\nu)$ is the eviatoric stress tensor with kinematic viscosity $\nu$, $\bm{g}$ represents the gravity, and $\kappa$ is the coefficient of molecular diffusivity. A key dimensionless parameter in the RT simulations is the Atwood number $A=(\rho_{h}-\rho_{l})/(\rho_{h}+\rho_{l})$, which quantifies the density contrast between the two fluids. For example, miscible fluids with greater density contrasts exhibit stronger instabilities. In this work, we adopt the dataset with $A=0.25$ provided in (Ohana et al., 2024) to train a solution operator $vec(\rho,\bm{u};t=t_{0})\mapsto vec(\rho,\bm{u};t=t_{0}+20)$ which evolves the system from rest states (with small perturbations) to mixed states after 20 time steps. The models are trained on 220 mapping pairs and tested on 40 pairs from distinct trajectories, using a downsampled grid of size $32\times 32\times 32$. Horizontal periodic boundary conditions and impermeable vertical walls are imposed. See Tab.2 for results.
@@
    To compare the baseline models in a global context of spatio-temporal dynamics, we evaluate their performance on predicting the next-hour ERA5 global wind field (Hersbach et al., 2020). All models are trained using data from the previous four days to predict the next-hour wind field at a grid size of $512\times 512$, and are tested on data from a new day with the same grid size. See Tab.15 for results.
We emphasize the broader learning difficulty spectrum of the benchmarks we selected. As shown in Fig. 5 (Li et al., 2021) which illustrates the spectral decay of truncation modes, our chosen benchmarks 6 and 7 contain significantly more high-frequency modes than simpler benchmarks such as 3 or other incompressible NS benchmarks with $\nu>1\times 10^{-4}$. To further evaluate the capability of these efficient and lightweight neural operators in capturing complex non-linear dynamics in 3D space, we include two additional, extremely challenging benchmarks: 8 and 9. Moreover, we also compare the performance of all baseline models in the rapid prediction of global wind fields, providing a comprehensive evaluation across both highly nonlinear and large-scale spatio-temporal dynamics.
@@
  • Mainly two types of Positional Features (Pos. Feat.), absolute coordinates (Abs.) and reference grid (Ref.), are concatenated into input signals: (1) FNO (Li et al., 2021) concatenates the input function value $a\in\mathbb{R}^{d_{a}}$ with the normalized absolute position coordinate $x\in\mathbb{R}^{d}$, usually scaled between 0 and 1. (2) Transolver (Wu et al., 2024a), on the other hand, uses an advanced version which constructs a coarse reference grid and computes Euclidean distances from each input point to each point in this reference grid, concatenating these relative positions $x_{ref}\in\mathbb{R}^{(d_{ref})^{d}}$. In our work, we utilize the same absolute positional features as FNO to ensure a fair comparison.
@@
  • For Transolver, the optimizers and scheduler type is AdamW (Loshchilov and Hutter, 2019) and OneCycleLR, while for FNO and SKNO are Adam (Kingma and Lei Ba, 2015) and Cosine Annealing. All the learning rates are initially set to $0.001$. More configurations can be found in the code provided in our supplementary materials.
@@
Averaging Neural Operator on $D_{p}$
In experiments governed by non-linear equations, we utilize ${F}_{p}^{-1}\tilde{A}{F}_{p}+B$, where $\tilde{A}$ and $B$ are both matrices with diagonally enhanced initialization. Such kind of implementation can be thought as an Averaging Neural Operator (ANO) (Lanthaler et al., 2025) along $p$, which governs the universal approximation of the PDE operator on $D_{p}$.
@@
Appendix C Related Work
@@
Unlike traditional methods (Lucia et al., 2004; Chinesta et al., 2011) that rely on dimension reduction techniques, such as symmetry analysis (Cantwell, 2002), to represent equations within their original domain, approaches like the level set method (Osher and Fedkiw, 2001; Osher et al., 2004; Adalsteinsson and Sethian, 1995; Jin and Osher, 2003) and the recent Schrödingerisation method (Jin and Liu, 2024; Jin et al., 2023, 2024) propose extending original PDEs to one more higher dimensional domains to create a pipe for functions to evolve. Similar "adding one dimension" ideas have also been used in signal processing like wavelets (Meyer, 1992) which employ scale and shift dimensions to represent original signals, and in quantum mechanics like the Wigner functions (Wigner, 1932) which utilize a single function displaying the probability (quasi-)distribution in the phase space.
@@
In scenarios where PDEs are unknown or training data is limited, supervised learning methods are employed to approximate solution operators, which recently shows their capacity to universally represent solution operators for partial differential equations (PDEs) (Hornik et al., 1989; Kovachki et al., 2021; Lu et al., 2021; Lagaris et al., 1998). Existing approaches mainly fall into two paradigms: Common Neural Networks and Neural Operators.
@@
Common NNs approximate solution operators working on a fixed grid size, which maps discretized input to output signals using architectures such as MLPs (He et al., 2016; Chen et al., 2018), CNNs (Qu et al., 2022; Brandstetter et al., 2023). For example, (Obiols-Sales et al., 2020; Khoo et al., 2021) utilizes a typical CNN structure to map the input to the output image, which is composed of the solution of numerical solvers.
@@
Neural operators (NOs) are designed to map between function spaces, enabling generalization across varying input discretizations without retraining (Kovachki et al., 2023; Wang et al., 2023). One of the most popular NO definitions is the kernel integral operator (Li et al., 2021; Raonic et al., 2023b; Cao et al., 2024). Early kernel function implementations employed graph-based message-passing networks (Li et al., 2020a, b), where dense edge connections encoded these kernels. However, these methods faced scalability challenges due to long training times and inefficiencies in modeling interactions across multiple spatial and temporal scales. Transformer-based models (Kissas et al., 2022), similar to graphs, also encounter limitations in computational efficiency due to quadratic time complexity. To address these challenges, tricky computation strategies on kernels have been proposed. Galerkin-based methods (Cao, 2021; Ern and Guermond, 2004) employ linear projections to parameterize kernels, while (Xiao et al., 2024) introduces a decomposition constrained by positive semi-definite assumptions. However, such methods suffer from the problems of reduced expressiveness due to the absence of non-linear activations like softmax. Recent SOTA model transolver (Wu et al., 2024a) proposes a physics-attention module to approximate the kernel function with the aid of a self-attention-based aggregation on flattened data. While these graph or transformer-based advanced methods try to escape from the heavy quadratic complexity by tricky implementations on the kernel functions originally defined in (Li et al., 2020a, 2021), they still pay the computational price for their neglect of the framework and model alignment in embedding spaces. While Mamba-based methods such as (Tiwari et al., 2025) can further reduce the computational cost over the domain after aggregation, the representational and complexity bottlenecks in the embeddings remain unaddressed. Instead, empirically designed fully connected or multi-head architectures are typically retained by default to model the evolution along the auxiliary dimension.
@@
Additionally, Fourier Neural Operators (FNO) (Li et al., 2021) efficiently implement spectral kernel representations by global convolutions via the Fast Fourier Transform (FFT). While training fast, it also faces the representation bottleneck like other NO methods because of the unaligned evolution in embedding spaces. Subsequent works take efforts to extend the Fourier Transform to other analytically well-explored transformations. For example, Complex Neural Operator (Tiwari et al., 2023) generalizes the Fourier Transformation with an extra learnable rotation parameter to model the $d$-dimensional kernel integration.
@@
There is another NO definition: DeepONet (Lu et al., 2021) employs a branch-trunk architecture inspired by the universal operator approximation theorem (Chen and Chen, 1995). These representations are combined through a dot product for the embeddings composed with functional outputs to predict the solution values. Recent advancements, such as integrating numerical solvers with DeepONet (Zhang et al., 2024), have improved convergence by capturing multi-frequency components, enhancing performance on complex PDE systems. However, its general-purpose design may limit its effectiveness in scenarios requiring high accuracy and PDE-aligned structural representations.
Appendix D Properties of SKNO
We highlight several properties of SKNO below.
Complexity
@@ Appendix D Properties of SKNO
To analyze computational complexity across general dimensional settings (1D, 2D, 3D, …), let $N_{x}$ denote the number of grid points in the spatial domain, $N_{p}$ the grid size along the auxiliary dimension, and $k_{x}$ the number of aggregated slices/nodes (used in Transolver). The complexity comparison is summarized in Tab. 8.
Table 8: Complexity comparison of different kernel implementations.
@@
Transolver $\mathcal{O}\!\left[N_{p}(N_{x}(k_{x}+N_{p})+k_{x}^{2})\right]\;(k_{x}\geq 32,\;N_{p}=128\text{ or }256)$
FNO $\mathcal{O}\!\left[N_{p}(N_{x}\log(N_{x})+N_{p}k_{x})\right]\;(k_{x}\leq 16,N_{p}\leq 32)$
SKNO (Ours) $\mathcal{O}\!\left[N_{p}(N_{x}(\log N_{x}+\log N_{p})+N_{p}k_{x})\right]\;(k_{x}\leq 16,N_{p}\leq 32)$
Table 9: Complexity comparison of different lifting implementations with positional encoding. (Assume all choose one linear layer)
Models Complexity
Transolver $\mathcal{O}\!\left[N_{p}(d_{a}+(d_{ref})^{d})\right]\;(d_{ref}=8,\;N_{p}=128\text{ or }256)$
FNO $\mathcal{O}\!\left[N_{p}(d_{a}+d)\right]\;(N_{p}\leq 32)$
SKNO (Ours) $\mathcal{O}\!\left[N_{p}(d_{a}+d)\right]\;(N_{p}\leq 32)$
Consider an extreme but practically relevant case in high-accuracy numerical simulations: a 3D field with grid resolution $((2)^{9})^{3}=(512)^{3}$ and $N_{p}=32$ (while Transolver typically requires $N_{p}=128$ for stable performance). Even with the minimal $k_{x}=32$, Transolver’s kernel implementation with a large linear constant ($k_{x}+N_{p}>128$) leads to a computational cost much higher than SKNO ($\log N_{x}=27$) or FNO ($\log N_{x}N_{p}=27+5=32$).
In lower dimensions like 1D or 2D cases, where $N_{x}$ is exponentially smaller, our kernel retains the linear-complexity advantages while achieving superior performance with significantly smaller scale constants compared to the SOTA Transolver (under the log benefit from Fast Fourier Transform). Here, we report the FLOPs of the kernel integration on representative experiments in Tab. 10. Notably, when the positional encodings are removed in Table 9, the prediction error of Transolver increases dramatically (from 6e36e-3 to 6e16e-1 in relative $L_{2}$ error on Burgers), whereas the other two models maintain nearly the same accuracy level.
Table 10: Computational cost comparison (FLOPs) of the kernel integration and relative rate (normalized by FNO).
Model Metric FNO Transolver SKNO
2D Navier–Stokes
FLOPs $16\times 10^{8} 1.1$$19\times 10^{9}$ $2.12\times 10^{8}$
Rate $1.000$ $10.292$ $1.831$
3D Rayleigh–Taylor Instability
FLOPs $8.98\times 10^{8}$ $9.00\times 10^{9}$ $1.59\times 10^{9}$
Rate $1.000$ $10.029$ $1.774$
Spectral Operator along $p$
Since our SKNO adopts a spectral implementation, we primarily compare it to the FNO (Li et al., 2021) as a reference framework. This choice also facilitates generalization to other spectral models like (Guibas et al., 2022; Li et al., 2024). To avoid notational clutter, we begin with the $d=1$ case, noting that the following arguments extend straightforwardly to higher dimensions.
Theorem D.1 (Neural Operator along the Auxiliary Dimension).
@@
$\hat{\bm{v}}_{i}:=F^{x}_{k_{x}}[P^{x}_{k_{x}}v](\hat{\bm{x}}_{i},\bm{p})\in\hat{V}_{i}\subset\mathbb{C}^{N_{p}},\qquad\tilde{\hat{\bm{v}}}_{i}:=F^{p}[R^{p}_{k_{p}}(F^{x}_{k_{x}}[P^{x}_{k_{x}}v]](\hat{\bm{x}}_{i},\tilde{\bm{p}}))\in\tilde{\hat{V}}_{i}\subset\mathbb{C}^{N_{p}},$
@@
    Approximating an operator $\mathcal{M}$ is equivalent (via the Fourier-conjugate decomposition) to approximating the conjugate mapping(s) $\{\widehat{\mathcal{M}}_{i}\}_{i=1}^{k_{x}}$ (and, after also Fourier-transforming in $p$, the conjugates $\{\widetilde{\widehat{\mathcal{M}}}_{i}\}$).
@@
    The forward DFT/frequency-projection and the inverse discrete transform admit neural operator approximation on the spectral domain (See Lemma 7 and Lemma 8 in (Kovachki et al., 2021), hence the operator learning problem reduces to a finite-dimensional neural-network approximation on the compact coefficient sets $\hat{V}_{i},\ \tilde{\hat{V}}_{i}$.
@@
Note. $\widehat{\mathcal{M}}_{i}:\hat{V}_{i}\subset\mathbb{R}^{2N_{p}}\rightarrow\mathbb{R}^{2N_{p}}$ for $i=1,...,k_{x}$ and $\widetilde{\widehat{\mathcal{M}}}_{i}:\tilde{\hat{V}}_{i}\subset\mathbb{R}^{2N_{p}}\rightarrow\mathbb{R}^{2N_{p}}$. Usually, we take $k_{p}=N_{p}$ and the $\frac{length(\bm{p})-k_{p}}{length(\bm{p})}$ random dropout $R^{p}_{k_{p}}=I^{p}$.
Proof. By the Fourier-conjugate decomposition (Eq. (12) in (Kovachki et al., 2021)), the operator can be written as
$\mathcal{M}$ $=\;(F^{x}_{k_{x}})^{-1}\circ\widehat{\mathcal{M}}_{k_{x}}\circ F^{x}_{k_{x}}\circ P^{x}_{k_{x}}$
$=\;(F^{x}_{k_{x}})^{-1}\circ(F^{p})^{-1}\circ\widetilde{\widehat{\mathcal{M}}}_{k_{x}}\circ F^{p}\circ F^{x}_{k_{x}}\circ P^{x}_{k_{x}}.$ (33)
@@
The mapping in Equ.33 $F^{p}F^{x}_{k_{x}}[P^{x}_{k_{x}}(\cdot)]$ sends the compact input family $V$ into a compact subset of Fourier coefficient space. Lemma 7 and Lemma 8 of (Kovachki et al., 2021) guarantee that the left-hand and right-hand parts of both $\widehat{\mathcal{M}}_{k_{x}}$ and $\widetilde{\widehat{\mathcal{M}}}_{k_{x}}$ can be approximated within the desired accuracy.
@@
Thus, the problem reduces to approximating the conjugate operators $\widehat{\mathcal{M}}_{k_{x}}$ or $\widetilde{\widehat{\mathcal{M}}}_{k_{x}}$, both of which are continuous maps defined on compact subsets of $\mathbb{R}^{2N_{p}}$. By the universal approximation theorem for neural networks (Barron, 2002; Hornik et al., 1989), any such continuous map can be approximated to arbitrary accuracy by a finite neural network. Hence, there exists finite NNs that approximate the conjugate mappings on $\hat{V}_{i}$ (or $\tilde{\hat{V}}_{i}$) within the desired accuracy.